# Blockchain Amplification Attack

### v2

@@ 6.2 Ratio of modified nodes: $\gamma$
Based on the HEAD of the git commit in the node name (which is generated when compiling the source code), we determine whether the node uses the publicly available version of the software or the private forks of the repository. 174 out of 175 peers share an identical git commit hash, which is not present in the corresponding GitHub repository. (We fail to collect information for one remaining node.) This fact ensures that an invalid transaction reaches out to all modified nodes. The details of how we extract the git commit hash and use it to determine whether a node was customized can be found in Appendix E. Further, we find that these modified nodes apply an aggressive propagation strategy ($\pi$=“aggressive”) – i.e., they only broadcast (0x02), but never announce (0x08), further increasing amplification effects.
Figure 8: Ratio of modified peers in our nodes’ connections.
6.3 Amplification factor
Based on the estimated parameters in the above sections, we finally calculate the amplification factors. We derive TAF based on the transaction duplication case, but the same technique could apply to insufficient balance or past nonce cases. Table II summarizes the model parameters and the final results. From the top, we refer to the size of transaction $a=560$ (bytes) in §5, modified node ratio $\gamma=0.015$, and propagation policy $\pi$=“aggresssive” in §6.2. We set the number of peers $N$ to 6,000 based on the statistics reported by [18, 17] after excluding non-Ethereum nodes. We plug in those numbers to Eqn. (2) and estimate the total waste in the network to be 2,037,613 bytes (2.0 MB). Given that the transaction size ($a$) is 560, the amplification factor is 3,638.
TABLE II: Estimated parameters for the model.
@@ 6.3 Amplification factor
N Number of nodes 6000 [18, 17]
@@
To quantify the economic impact of network waste (EAF), we next calculate the financial loss of spreading invalid transactions by modified nodes. We observe that many centralized services deploy their nodes in the cloud service (Appendix F), so we calculate the data transfer cost based on the “pay-as-you-go” policy of the cloud services. We choose AWS pricing [^8] https://aws.amazon.com/ec2/pricing/on-demand/#Data_Transfer_within_the_same_AWS_Region as an example. AWS does not charge for inbound traffic but for outbound traffic (plus, significantly less for the traffic within AWS). Most cloud services follow a similar standard. If we assume that the cost of inbound traffic is zero, the attack incurs costs only for 1) the attacker (to send invalid transactions) and 2) the modified nodes (to propagate invalid transactions to the rest).
@@
We compute EAF (economic amplification factor) based on TAF and the traffic price ratio, given by Eqn. (1). If the attacker strategically deploys its node in the cloud service co-located with modified nodes, the cost is 0–20 USD/TB (AWS US East) including the case when the attacker sends transactions to multiple modified nodes across regions. On the one hand, the modified nodes are organically connected to the rest of the P2P network, with 20% of internal AWS traffic (Appendix F) and 80% of external traffic to the internet. AWS dynamically charges external traffic starting from 90 USD/TB up to 10TB; the price gets discounted after.99990 USD/TB until 10TB, 85 USD/TB for the next 40TB, 70 USD/TB for the next 100TB, and 50 USD greater than 150TB Due to the price discrepancies between the attacker and the modified nodes (i.e., price multiplier), EAF is 13,827 for the first 10GB of non-AWS traffic (12.5 GB for all). Figure 9 illustrates the decrease in EAF ($y$-axis) based on the amount of outgoing traffic ($x$-axis) due to AWS’s dynamic pricing. The attacker’s economic benefit marginally decreases over the scale of the attack.
Figure 9: EAF based on the traffic usage of a modified node.
7 Attack simulation in the local P2P network
@@
We evaluate our proposed attack in the local P2P network simulation and compare it with the existing DoS attacks.
@@
We set up a local P2P network, in which we deploy our modified node, attack the node, and then assess both txpool congestion and transaction inclusion in the blocks. We base our attack on Yaish et al.’s publicly available repository, [^10] https://github.com/AvivYaish/SpeculativeDoS which allows for a direct comparison of our approach with the two existing DoS attacks, namely 1) Baseline, that is, sending valid transactions with higher gas prices (naive eviction strategy), and 2) MemPurge [55], that is, sending future latent (invalid) transactions by extending the DETER method [36]. We test all three attacks in our modified node.
@@
We set up one validator, 80 honest accounts, and a varying number $x$ of malicious accounts, all of which interact with our node through the RPC endpoint. First, each honest account sends 64 transactions to collectively fill up the transaction pool with valid transactions. Each attacker then begins submitting attack transactions. For our proposed attack, each attack account attempts to transfer an amount higher than its current account balance, and starting from nonce 0, which is lower than the current nonce – i.e., these transactions are invalid. Each attack account then gradually increments the nonce to generate 32 invalid transactions (i.e., a total of $32x$ attack transactions) with the intention of evicting honest transactions.
@@
Figures 10 (middle and right) show that our proposed attack can exclude honest transactions not just in the txpool but block as well because the attacker’s invalid transactions occupy the txpool. Furthermore, these invalid transactions are never included in the block, resulting in zero attack transaction costs. MemPurge gradually displaces honest transactions in the block, while its attacker pays for one transaction per attack account. The Baseline attack replaces honest transactions with a smaller number of attack accounts. Yet, all the attack transactions are included in the block, resulting in a significantly high(er) attack cost.
In summary, our attack achieves the same level of txpool/block eviction rate as the two existing DoS attacks but with significantly lower (zero) attack transaction costs. This advantage allows the attacker to keep refreshing the txpool with its new invalid transactions, causing the modified node to continuously notify its neighbors, thereby incurring egress traffic loss—EDoS attack. In Appendix G, we describe more detailed attack scenarios and roughly estimate the maximum traffic monthly costs incurred at the modified node to be 88,904 USD per node (8M USD in aggregate) when an attacker saturates outgoing traffic with invalid transactions.
We open-source our attack code in our repository111111hhttps://github.com/taro-tsuchiya/BlockchainAmplification and document the attack implementation details.
Figure 10: The ratio of honest transactions in the txpool (left), the ratio of honest transactions in the block (middle), the number of attack transactions in the block (right), over the number of attack accounts (x-axis).
8 Estimating transaction validations
In this section, we conduct two experiments: 1) simulation, and 2) empirical measurement, focused on the latency benefits to understand the economic rationale behind skipping certain transaction validation steps.
@@ 8 Estimating transaction validations
8.1 Simulating validations
We first create an isolated fork of the Ethereum blockchain within our local network to estimate the Geth validation processing time. We use a modified Geth client, integrated with the Pebble database. [^12] Geth can run on either Pebble or LevelDB as its underlying database. This approach allows us to simulate a controlled environment without impacting the live P2P network. Our objective is to examine how different types of validations contribute to the overall processing time and identify potential areas for efficiency improvements. Specifically, we run two functions in the first “accept” stage: ValidateTransaction() and ValidateTransactionWithState() for 1,000 consecutive blocks, starting from block number 18,140,000, in a total of 136,437 transactions. We conduct our experiment on a Ubuntu 22.04.2 LTS server, equipped with an AMD Ryzen Threadripper 3990X 64-Core Processor and a Corsair MP400 NVMe disk.
@@ 8.1 Simulating validations
8.2 Empirically measuring validations
The previous experiment solely examines validation processing time, omitting txpool or the timing of transaction arrival. We next empirically measure the latency difference between modified and regular nodes. Since our modified node may forward attackers’ invalid transactions, we choose to run the node in the Ethereum testnet “Holesky” instead of the mainnet. Our modified node bypasses the first (“accept”) and the second (“forward”) validation stages, but keeps the third (“remove”) validation stage, because the node gets congested by past transactions and stops inserting new transactions. To validate our implementation, we send one test invalid transaction (insufficient balance and past nonce) through our node’s RPC and verify that this transaction appears in our modified nodes’ txpool. To limit confounding factors, we alternate between running an unmodified and a modified node on the same machine. This allows us to control for 1) the specs of the machine (Ubuntu 22.04.4 LTS, with an AMD EPYC 9124 16-Core processor) and 2) the network topology by connecting to a similar set of peers using the same public key. To further eliminate the effect of peer connections, we increase the peer limit from the default (50) to 300 to connect to as many active peers as possible (based on the observation by Zhao et al [56] on the Goerli testnet). We sync the blockchain ahead of time and run our node from April 4, 2024, 1 PM (UTC) to July 12, 2024, 1 PM. We switch the modified node and the regular node weekly or bi-weekly basis, resulting in 7 weeks for each node. We capture around 34 million transactions in total.
To compare performance, we monitor each txpool and record the timestamp when the node finishes processing/validating each transaction. We use a metric called “inclusion time” [51]: the timestamp of blockchain minus the first observation timestamp in our txpool. The longer the inclusion time, the faster the node discovers and processes transactions. Figure 11 illustrates the distribution (until 120 seconds for visibility) for inclusion time for the modified (right) and the regular (left) node, respectively. We exclude periods when the market behaves abnormally (e.g., due to attacks or testing); refer to Appendix I.
@@ 8.2 Empirically measuring validations
Figure 11: Time inclusion for a modified/regular node
9 Discussion
We first discuss if, despite the potential for losses coming from attacks, it could still be economically rational for an operator to modify nodes. Second, we describe the limitations and the possible mitigations against the attack.
9.1 Cost-benefit analysis
We summarize the benefits and the costs for each player: an attacker, modified nodes, and regular nodes. The attacker can launch an attack to induce financial losses for modified nodes. For instance, an attacker economically competing with the entity operating modified nodes could use the attack to weaken the competition. The attacker can also use its invalid transactions to displace legitimate transactions, rendering the service unusable (i.e., causing a DoS) as shown in §7.
@@ 9.1 Cost-benefit analysis
Based on §7 and §8.1, while omitting validation checks leads to a possible reduction in processing time by the order of milliseconds, the marginal time savings may not necessarily justify the potential damages/risks from the Blockchain Amplification Attack (in terms of MEV profits). Modified nodes engaging in these “optimizations” also expect to face degradation in the quality of service provided, due to the increasing number of invalid transactions (as shown in §7). With this in mind, a rationally economic node would still continue skipping validations as long as the economic benefits (including revenue from users paying to connect to those services) outweigh any financial losses incurred from the attack.
@@
Regular nodes can expect to receive transactions faster by connecting to the modified nodes but receive more invalid (less valid) transactions, which consume network capacity and CPU, indicating that our attack poses a threat to the security of all the players in a P2P network. (Our attack does not cause congestion in these nodes’ txpool, since regular nodes discard invalid transactions.)
TABLE III: Estimation of profit increase vs. latency reduction per block in Ethereum based on Wahrstätter et al. [52]. Calculations assume an ETH price of \$2,500, and an average MEV bid of 0.06 ETH.
Submit the block at $0.409s$ $1s$ $2s$ $2.5s$
For Every 1ms Latency Reduction
% Profit Increase $0.034\%$ $0.031\%$ $0.018\%$ $0.0074\%$
9.2 Limitations and mitigations
A few key assumptions affect our estimation model. First, we assume that the set of nodes we are connected to is a representative sample of the overall population. However, if our node(s) is/are more likely to connect to the node with many open connections, we might overestimate the distribution of peer connections $g(x)$.
Second, our method of estimating the number of active peers requires the assumption of the network being static over time. We mitigate this concern by only including the nodes with a significant number of messages—stable nodes, thereby reinforcing the validity of our assumption.
Third, while our model assumes that the node transmits each invalid transaction once, in practice, the node can resend the same transaction multiple times, which could significantly increase the amplification factor.
@@ 9.2 Limitations and mitigations
Possible mitigations. Previous work suggests a few countermeasures to prevent DoS/EDoS attack [5]: 1) testing the client (Turing test, cryptographic puzzle) to exclude non-human requests or 2) blocklisting users by detecting abnormal traffic. Those solutions do not directly translate to our context because 1) a blockchain node is not interactive (i.e., a human would not be able to answer challenges at the requested rate) and 2) node identity can be easily spoofed. Unlike previous blockchain DoS attacks [36, 55, 28], our identifid vulnerability does not exist in the public client, thereby patching the current client does not resolve the issue. We devise three solutions tailored to our case.
@@
First, one can enforce a stricter txpool policy (checking balance, nonce, gas price bump) on modified nodes. Upon recovering more stringent transaction checks in the modified nodes, attack vectors diminish. However, as long as profits could potentially exceed losses on modified nodes, we cannot expect the solution to be universally implemented.
@@
Second, one could postpone the validation process by introducing a two-step propagation mechanism. The idea is similar to the solutions proposed by Das et al [11] for delaying block validation in the context of the Verifier’s Dilemma. After the modified node receives transactions, it 1) directly relays them to users without validation, ensuring latency reduction, 2) validates transactions afterward, and 3) sends them to the rest (or does not forward them at all). While this two-step approach does not negatively impacts the rest of the network, modified nodes still pay the cost of delivering invalid transactions to users, creating a risk for users to include invalid transactions.
@@
Third, one could employ a reputation system on regular nodes (e.g., similar to ISP blocklisting [6]). If the (regular) nodes have a mechanism to disconnect from the modified nodes that propagate many invalid transactions, the modified nodes would stop hearing new transactions, getting isolated from the rest of the network. This approach might yield false positives (i.e., legitimate nodes being blocked due to discrepancies in synced statuses) or lead to a cat-and-mouse game (i.e., adversarial nodes could come up with techniques to bypass/game the reputation system) [23].
10 Related work
We discuss related work by examining previous efforts in P2P network measurements and network security.
10.1 Measurement on P2P network
Most of the measurements on blockchain P2P networks focus on describing the topology of the network such as network size [31, 44], influential nodes [35, 38], stability or longevity of the connections [43, 10], and latency between peers [12, 21, 48, 51]. Some websites summarize those statistics in real-time for various networks (e.g., Bitcoin [2, 7, 47]; Ethereum [18, 17, 40]; Monero [42])
In particular, there has been extensive work on identifying active peer connections to further understand the propagation flow over the network. The techniques developed include: utilizing latency between peers [44, 10], isolation property and orphan transactions [13], connection freshness [4], and gas fee change [35, 56]. However, these works are either 1) chain-specific and do not apply to Ethereum, and/or 2) involve active measurement (e.g., submitting transactions), which is expensive and could disrupt the main network.
Our proposed method in §6.1 works on the Ethereum main network 1) without intervening in ongoing activities and 2) without incurring any additional costs apart from collecting data. Our method only considers the number of connections and does not determine whether an active edge exists between two specific nodes, making it harder to compare its accuracy/performance with previous proposals.
Another body of measurement research focuses on characterizing nodes, e.g., centralized entities such as mining or relay nodes [41, 10, 35, 36], AS/cloud services [20, 21, 44], geo-distributions [43, 31, 10, 38, 4, 8], and peer technical specifications (i.e., bandwidth, CPUs) [48, 8]. Closest to our work is the literature about the client (software) types – for Bitcoin [43], Ethereum [31, 8, 24], and Zcash [10]. Those works mostly focus on the diversity of software and/or versions, with relatively scarce attention on the implementation and performance. To the best of our knowledge, there has not been any exploration of customized clients, a gap our study aims to bridge.
10.2 P2P network security: DoS and spam
Due to the rise of centralized services such as mining pools, exchanges, and relay nodes, Denial of Service (DoS) attacks on those services can have a disastrous impact on the network [50]. The most directly relevant work is by Li et al [36] (and its extension by Yaish et al [55]), which exploits the txpool’s propagation policy by sending transactions with a future nonce to evict legitimate transactions in the txpool. Further, Wang et al [54] developed a fuzzer to automatically discover txpool vulnerabilities. Heo et al [28] illustrate that attackers can delay honest transaction execution by sending invalid transactions. Zhou et al [57] document the spamming behavior of MEV bots as a computational overhead to the network. Our paper describes a new Blockchain DoS attack, exploiting the lack of transaction validations—a superset of the above DoS-related works, as shown in §7), and empirically derives the amplification factors. We also reference some metrics from non-blockchain DoS literature such as amplified DoS and EDoS attacks from Kumar [33] and Wang et al. [53], respectively.
11 Conclusion
A reduction in latency creates an economic value as a form of MEV/BEV, and some entities modify the software client (“modified node”) to shorten the transaction validation process. We formalize the Blockchain Amplification Attack, identify similar attack instances in the wild and empirically measure and simulate the attack impact. Our model concludes that an attacker can amplify its original traffic by 3,600 times, and the financial loss by 13,800 times. A lack of transaction validation not only increases traffic costs on the modified nodes but also introduce invalid transactions, and degrades the user experience. Our modeling frameworks provide a foundation for explaining how attack transactions propagates through the network and pose a threat to the overall blockchain ecosystem.
12 Ethics

### v3

@@ 6.2. Ratio of modified nodes: $\gamma$
Based on the HEAD of the git commit in the node name (which is generated when compiling the source code), we determine whether the node uses the publicly available version of the software or the private forks of the repository. 174 out of 175 peers share an identical git commit hash, which is not present in the corresponding GitHub repository. (We fail to collect data for one remaining node.) This fact suggests that those nodes are likely from the same entity and ensures that an invalid transaction reaches out to all modified nodes. The details of how we extract the git commit hash and use it to determine whether a node was customized can be found in Appendix E. Further, we find that these modified nodes apply an aggressive propagation strategy ($\pi$=“aggressive”) – i.e., they only broadcast (0x02), but never announce (0x08), further increasing amplification effects.
Figure 7. Ratio of modified peers in our nodes’ connections.
@@
N Number of nodes 6000 (Etherscan.io, 2023; Ethernodes.org, 2023)
@@
Table 2. Estimated parameters for the model.
6.3. Amplification factor
Based on the estimated parameters in the above sections, we finally calculate the amplification factors. We derive TAF based on the transaction duplication case, but the same technique could apply to insufficient balance or past nonce cases. Table 2 summarizes the model parameters and the final results. From the top, we refer to the size of transaction $a=560$ (bytes) in §5.2, modified node ratio $\gamma=0.015$, and propagation policy $\pi$=“aggresssive” in §6.2. We set the number of peers $N$ to 6,000 based on the statistics reported by the third-party websites (Etherscan.io, 2023; Ethernodes.org, 2023) after excluding non-Ethereum nodes. All in all, we estimate the total waste per transaction in the network to be 2,037,613 bytes (2.0 MB). Given that the transaction size ($a$) is 560, the amplification factor is 3,638.
@@ 6.3. Amplification factor
To quantify the economic impact of network waste (EAF), we next calculate the financial loss of spreading invalid transactions at modified nodes. We observe that many centralized entities deploy their nodes in the cloud service (Appendix F), so we calculate the data transfer cost based on the “pay-as-you-go” policy of the cloud services. We choose AWS pricing [^8] https://aws.amazon.com/ec2/pricing/on-demand/#Data_Transfer_within_the_same_AWS_Region as an example. AWS does not charge for inbound traffic but for outbound traffic (plus, significantly less for the traffic within AWS). Most cloud services follow a similar standard. If we assume that the cost of inbound traffic is zero, the attack incurs costs only for 1) the attacker (to send invalid transactions) and 2) the modified nodes (to propagate invalid transactions to the rest).
@@
We compute EAF (economic amplification factor) based on TAF and the traffic price ratio, given by Eqn. (1). If the attacker strategically deploys its node in the cloud service co-located with modified nodes, the cost is 0–20 USD/TB (AWS US East) including the case when the attacker sends transactions to multiple modified nodes across regions. On the one hand, the modified nodes are organically connected to the rest of the P2P network, with 20% of internal AWS traffic (Appendix F) and 80% of external traffic to the internet. AWS dynamically charges external traffic starting from 90 USD/TB up to 10TB. EAF is 13,827 for the first 10GB of non-AWS traffic.
7. Attack simulation in the local P2P network
@@
To precision the attack mechanics, we simulate our proposed attack in the local network and compare it with the existing DoS attacks.
@@
We set up a local P2P network, in which we deploy our modified node, attack the node, and then assess both txpool congestion and transaction inclusion in the blocks. We base our attack on Yaish et al.’s publicly available repository, [^9] https://github.com/AvivYaish/SpeculativeDoS which allows for a direct comparison of our approach with the two existing DoS attacks, namely 1) Baseline, that is, sending valid transactions with higher gas prices (naive eviction strategy), and 2) MemPurge (Yaish et al., 2023), that is, sending future latent (invalid) transactions by extending the DETER method (Li et al., 2021b). We test all three attacks in our modified node.
@@
We set up one validator, 80 honest accounts, and a varying number $x$ of attack accounts, all of which interact with our node through the RPC endpoint. First, each honest account sends 64 transactions to collectively fill up the txpool with honest transactions. Each attacker then submit attack transactions to evict honest transactions. In our proposed attack, a attacker attempts to transfer an amount higher than its current balance, and starting from nonce 0, which is lower than the current nonce – i.e., these transactions are invalid. Each attack account then gradually increments the nonce to generate 32 invalid transactions (i.e., a total of $32x$ attack transactions).
@@
Figures 8 (middle and right) show that our proposed attack can exclude honest transactions not just in the txpool but block as well because the attacker’s invalid transactions occupy the txpool. The delay (or the failure) in the transaction inclusion negatively impacts the security of the other layers (e.g., consensus) (Gervais et al., 2015; Wahrstätter et al., 2024). Furthermore, these invalid transactions are never included in the block, resulting in zero attack transaction costs. MemPurge gradually displaces honest transactions in the block, while its attacker pays for one transaction per attack account. The Baseline attack replaces honest transactions with a smaller number of attack accounts. Yet, all the attack transactions are included in the block, resulting in a significantly high(er) attack cost.
In summary, our attack achieves the same level of txpool/block eviction rate as the two existing DoS attacks but with significantly lower (zero) attack transaction costs. This advantage allows the attacker to keep refreshing the txpool with its new invalid transactions, causing the modified node to continuously notify its neighbors, thereby incurring egress traffic loss—EDoS attack. In Appendix G, we describe more detailed attack scenarios and roughly estimate the maximum traffic monthly costs incurred at the modified node to be 88,904 USD per node (8M USD in aggregate) when an attacker successfully saturates outgoing traffic with invalid transactions.
We open-source our attack code in our repository [^10] https://github.com/taro-tsuchiya/BlockchainAmplification and document the attack implementation details.
Figure 8. The ratio of honest transactions in the txpool (left), the ratio of honest transactions in the block (middle), the number of attack transactions in the block (right), over the number of attack accounts (x-axis).
8. Estimating transaction validations
In this section, we conduct two experiments: 1) simulation, and 2) empirical measurement, focused on the latency benefits behind skipping certain transaction validation steps.
@@ 8. Estimating transaction validations
8.1. Simulating validations
We first create an isolated fork of the Ethereum blockchain within our local network (without impacting the live P2P network) to estimate the Geth validation processing time. We use a customized Geth client, integrated with the Pebble database. [^11] Geth can run on either Pebble or LevelDB as its underlying database. Our objective is to examine how different types of validations contribute to the overall processing time and identify potential areas for efficiency improvements. Specifically, we run two functions in the first “accept” stage: ValidateTransaction() and ValidateTransactionWithState() for 1,000 consecutive blocks, starting from block number 18,140,000, in a total of 136,437 transactions. We conduct our experiment on a Ubuntu 22.04.2 LTS server, equipped with an AMD Ryzen Threadripper 3990X 64-Core Processor and a Corsair MP400 NVMe disk.
@@ 8.1. Simulating validations
8.2. Empirically measuring validations
The previous experiment solely examines validation processing time, omitting txpool or the timing of transaction arrival. We next empirically measure the latency difference between modified and regular nodes. Since our modified node may forward attackers’ invalid transactions, we choose to run the node in the Ethereum testnet “Holesky” instead of the mainnet. Our modified node bypasses the first (“accept”) and the second (“forward”) validation stages, but keeps the third (“remove”) validation stage, because the node gets congested by past transactions and stops inserting new transactions. To validate our implementation, we send one test invalid transaction (insufficient balance and past nonce) through our node’s RPC and verify that this transaction appears in our modified nodes’ txpool. To limit confounding factors, we alternate between running an unmodified and a modified node on the same machine. This allows us to control for 1) the specs of the machine (Ubuntu 22.04.4 LTS, with an AMD EPYC 9124 16-Core processor) and 2) the network topology by connecting to a similar set of peers using the same public key. To further eliminate the effect of peer connections, we increase the peer limit from the default (50) to 300 to connect to as many active peers as possible (based on the observation by Zhao et al (Zhao et al., 2024) on the Goerli testnet). We sync the blockchain ahead of time and run our node from April 4, 2024, 1 PM (UTC) to July 12, 2024, 1 PM. We switch the modified node and the regular node weekly or bi-weekly basis, resulting in 7 weeks for each node. We capture around 34 million transactions in total.
To compare performance, we monitor each txpool and record the timestamp when the node finishes processing each transaction. We use a metric called “inclusion time” (Wahrstätter et al., 2024): the timestamp of blockchain minus the first observation timestamp in our txpool. The longer the inclusion time, the faster the node discovers and processes transactions. We exclude periods when the market behaves abnormally (e.g., due to attacks or testing); refer to Appendix I.
@@ 8.2. Empirically measuring validations
9. Discussion
We first discuss if, despite the potential attack damages, it could still be economically rational for an operator to modify nodes. We then describe the limitations and the mitigations against the attack.
9.1. Cost-benefit analysis
We summarize the benefits and the costs for each player: an attacker, modified nodes, and regular nodes. The attacker can induce financial losses for modified nodes. For instance, an attacker economically competing with the entity operating modified nodes could use the attack to weaken the competition. The attacker can also use its invalid transactions to displace legitimate transactions, rendering the service unusable (i.e., causing a DoS) as shown in §7.
@@ 9.1. Cost-benefit analysis
While omitting validation checks leads to a possible reduction in processing time by the order of milliseconds (§8.1), the marginal time savings may not necessarily justify the potential damages from the blockchain amplification attack (i.e., up to \$89,000 per node in §7). Modified nodes also expect to face degradation in the quality of service provided, due to the increasing number of invalid transactions (as shown in §7). With this in mind, a rationally economic node would still continue skipping validations as long as the economic benefits (including revenue from users paying to connect to those services as discussed in §2) outweigh any financial losses incurred from the attack.
@@
Regular nodes can expect to receive transactions faster by connecting to the modified nodes but receive more invalid (less valid) transactions, which consume network capacity and CPU. This fact indicates that our attack poses a threat to the security of all the players in a P2P network. (Our attack does not cause congestion in these nodes’ txpool, since regular nodes discard invalid transactions.)
9.2. Limitations and mitigations
A few key assumptions affect our estimation model. First, we assume that the set of nodes we are connected to is a representative sample of the overall population. However, if our node(s) is/are more likely to connect to the node with many open connections, we might overestimate the distribution of peer connections $g(x)$. Second, our method of estimating the number of active peers requires the assumption of the network being static over time. We mitigate this concern by only including the nodes with a significant number of messages—stable nodes, reinforcing the validity of our assumption. Third, while our model assumes that the node transmits each invalid transaction once, in practice, the node can resend the same transaction multiple times, which could significantly increase the amplification factor.
@@ 9.2. Limitations and mitigations
Possible mitigations. Previous work suggests a few countermeasures to prevent DoS/EDoS attack (Chowdhury et al., 2017): 1) testing the client (Turing test, cryptographic puzzle) to exclude non-human requests or 2) blacklisting users on abnormal traffic. Those solutions do not directly translate to our context because 1) a blockchain node is not interactive (i.e., a human would not be able to answer challenges at the requested rate) and 2) node identity can be easily spoofed. Unlike previous blockchain DoS attacks (Li et al., 2021b; Yaish et al., 2023; Heo et al., 2023), our identified vulnerability does not exist in the public client, thereby patching the current client does not resolve the issue. We devise three solutions tailored to our case.
@@
First, one can enforce a stricter txpool policy (checking balance, nonce, gas price bump) on modified nodes. Upon recovering more stringent transaction checks in the modified nodes, attack vectors diminish. However, as long as profits could potentially exceed losses on modified nodes, we cannot expect the solution to be implemented.
@@
Second, modified nodes could apply validations after they forward transactions to users. The idea is similar to the solutions proposed by Das et al (Das et al., 2021) for delaying block validation in the context of the Verifier’s Dilemma. After the modified node receives transactions, it 1) directly relays them to users (privately) without validation, ensuring latency reduction, 2) validates transactions afterward, and 3) sends them to the rest (or does not forward them at all). While this two-step approach does not negatively impacts the rest of the network, modified nodes still pay the cost of delivering invalid transactions to users, creating a risk for users to include invalid transactions.
@@
Third, one could employ a reputation system on regular nodes (e.g.,the ISP blacklisting (Clayton, 2009) or the penalty mechanism in the Bitcoin P2P network (Gervais et al., 2015)). If the (regular) nodes have a mechanism to disconnect from the modified nodes that propagate many invalid transactions, the modified nodes would stop hearing new transactions, getting isolated from the rest of the network. However, the Ethereum community is generally hesitant to introduce the node reputation system for two reasons. First, the system could produce false positives due to discrepancies in the synced status (i.e., invalid transactions being considered valid if the node lags behind) or differences in client implementations, leading to a partitioned network. Second, the system invites a cat-and-mouse game with adversaries that attempt to bypass/game the reputation system. See more discussions in the Geth community (Geth, 2024a, b)
9.3. Ethics